1.The product of two numbers is 8000. If LCM of two numbers is 800, find HCF?
2. Find the greatest number, which divides 149 and 101 leaving remainder 5 in each case?


Answer by ortarico(23)

Although your question should not have been posted on this section, here we go:
1. Let the two numbers be x and y
Now, if xy is 8000 and their L.C.M is 800. Therefore, the H.C.F of x and y = 8000/800
=> 10
2. 149 - 5 = 144
101 - 5 = 96
Therefore, find the H.C.F of 144 and 96. The answer is 48.
Hope the solutions helped?